# Xingjian Wu, Jianxin Jin, Wanghui Qiu, Peng Chen, Yang Shu, Bin Yang, Chenjuan Guo

### v2

@@ Abstract
Cross-domain generalization is very important in Time Series Forecasting because similar historical information may lead to distinct future trends due to the domain-specific characteristics. Recent works focus on building unimodal time series foundation models and end-to-end multimodal supervised models. Since domain-specific knowledge is often contained in modalities like texts, the former lacks the explicit utilization of them, thus hindering the performance. The latter is tailored for end-to-end scenarios and does not support zero-shot inference for cross-domain scenarios. In this work, we introduce Aurora, a Multimodal Time Series Foundation Model, which supports multimodal inputs and zero-shot inference. Pretrained on Corss-domain Multimodal Time Series Corpus, Aurora can adaptively extract and focus on key domain knowledge contained in corrsponding text or image modalities, thus possessing strong Cross-domain generalization capability. Through tokenization, encoding, and distillation, Aurora can extract multimodal domain knowledge as guidance and then utilizes a Modality-Guided Multi-head Self-Attention to inject them into the modeling of temporal representations. In the decoding phase, the multimodal representations are used to generate the conditions and prototypes of future tokens, contributing to a novel Prototype-Guided Flow Matching for generative probabilistic forecasting. Comprehensive experiments on well-recognized benchmarks, including TimeMMD, TSFM-Bench and ProbTS, demonstrate the consistent state-of-the-art performance of Aurora on both unimodal and multimodal scenarios.
@@ 1 Introduction
Time series forecasting has gained sustained attention for decades of years due to its significant values in multiple domains, including economy, transportation, meteorology, and public health. In recent years, the key pivot comes with the surge of deep learning, which brings the boom of merticulously-designed deep forecasting models (Cirstea et al., 2022; Nie et al., 2023; Qiu et al., 2025d; Wu et al., 2025a). Through learning the inherent dynamics within the raw data, deep learning models can outperform classic statistical methods (Box & Pierce, 1970; Mei et al., 2014) and obey the scaling law (Shi et al., 2024a; Yao et al., ). Due to the success, it also brings the most commonly-used forecasting paradigm, which utilizes the past information to infer how the series goes in the coming horizon. Although this paradigm contributes to impressive performance under the domain-specific scenarios, its effectiveness is suspicious when facing cross-domain inference, where similar historical information may lead to different futures due to domain differences.
As shown in Figure 1, current research of time series forecasting explores the cross-domain adaption problem in two main perspectives: 1) pre-training on cross-domain time series corpus for unimodal time series foundation models, which partially possess cross-domain generalization capablities; 2) utilizing cross-modality information in training end-to-end multimodal supervised models, which effectively integrates domain knowledge in forecasting. For time series foundation models, the cross-domain generalization capabilities mainly come from the sensitivity to subtle differences in historical information from different domains. Some of them (Shi et al., 2024b; Liu et al., 2025d) are pretrained on trillion-scale corpus with hevay backbones, thus possessing certain cross-domain adaption capabilities. Others (Wang et al., 2025c; Woo et al., 2024) have specific structures,
@@
which excels at capturing cross-domain features. However, their capabilites come from single time modality and lack explict domain knowledge guidance, thus hindering the performance. For end-to-end multimodal supervised models (Jin et al., 2024; Liu et al., 2025a), though they consider the multimodal knowledge to enhance the domain-specific forecasting, they lack the ability to support zero-shot forecasting in cross-domain scenarios. In our view, the aurora of next-generation time series foundation model lies in pretraining a cross-modality model on cross-domain time series corpus, which can utilize the domain knowledge within modalities and serve as a versatile out-of-box forecaster in complex scenarios.
@@
  • Experimentally, Aurora achieves state-of-the-art performance on three well-recognized benchmarks, including datasets from TimeMMD (Liu et al., 2024b), TSFM-Bench (Li et al., 2025c) and ProbTS (Zhang et al., 2024), covering both unimodal and multimodal scenarios, thus demonstrating a strong out-of-box tool of decision intelligence.
@@ 2.2 Time Series Foundation Models
To support cross-domain generalization, unimodal Time Series Foundation Models are widely studied. The majority of them adopt Tranformer-based architectures, which are pretrained on time series corpus of billion- or trillion- scale to obtain the strong generalization capabilities. Among them, Sundial (Liu et al., 2025d), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., ), UniTS (Gao et al., 2024), Timer (Liu et al., 2024e), and TimesFM (Das et al., 2024) demonstrate strong zero-shot forecasting performance on unimodal tasks, even outperforming those full-shot supervised models in many cases. Considering the forecasting paradigm, Sundial, MOIRAI, Chronos, and Lag-Llama (Rasul et al., 2023) also support probabilistic forecasting, which provides additional robustness and versatility for decision-making. Despite their endeavors to enhance cross-domain generalization capabilities, when historical series exhibit similarities, the forecasts they generate remain static. This lack of adaptability renders them unable to accommodate the diverse and changing real-world domains.
@@ 3.2.2 Prototype-Guided Flow Matching
Different from DDPM (Ho et al., 2020), which can be treated as an SDE solver to transform data from fixed Gaussian distributions to realistic target distributions, Flow Matching (Lipman et al., 2023) serves as a more intuitive and smooth ODE solver, which learns the Velocity Field between a random initial distribution and the target distribution. However, current methods (Liu et al., 2025d; Kollovieh et al., ) still set the initial distributions as Standard Gaussian, which neglects the capability of Flow Matching to work like a stochastic interpolant. Obviously, constructing a propriate prototype as the inital starting point can enhance the intuitiveness and stability of Flow Matching.
In the inference phase, the sampling process is a discretized integration process–see Algorithm 1. The gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,\mathbf{I})$ helps support probabilistic forecasting. Finally, we can obtain the forecasts $\hat{y}_{i}\in\mathbb{R}^{p^{time}}$ of the $i$-th future token. And the forecasts of the future horizon are $\hat{Y}=\text{Concat}\{\hat{y}_{i}\}\in\mathbb{R}^{F\times p^{time}}$.
@@ 4 Experiments
We make extensive experiments to evaluate the performance of Aurora. Specifically, we introduce the experimental settings in Section 4.1. In Section 4.2, we evaluate the zero-shot and few-shot performance of Aurora on multimodal forecasting scenarios. Considering the modal absence in the realistic world, we also evaluate the zero-shot performance of Aurora on unimodal forecasting scenarios–see Section 4.3. To analyze the key components in Aurora, we also make detailed model analyses in Section 4.4. In summary–see Figure 4, our proposed Aurora achieves state-of-the-art performance in both unimodal and multimodal forecasting scenarios.
@@ 4.1 Experimental Settings
Cross-Domain Multimodal Time Series Corpus. We first collect a substantial number of open-source time series datasets across diverse domains, then generate the corresponding sample-wise textual descriptions using Large Language Model (Liu et al., 2024a), which simulates the downstream scenarios with domain-specific textual information. The detailed method is introduced in Appendix A.
Benchmarks. We evaluate both the multimodal forecasting and unimodal forecasting performance of Aurora on well-recognized benchmarks, including TimeMMD (Liu et al., 2024b), TSFM-Bench (Li et al., 2025c), and ProbTS (Zhang et al., 2024). Note that these benchmarking datasets are strictly excluded from the pretraining time series corpus.
Baselines. We compare Aurora with 10 well-known unimodal time series foundation models, including Sundial (Liu et al., 2025d), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TimesFM (Das et al., 2024), Timer (Liu et al., 2024e), UniTS (Gao et al., 2024), Chronos (Ansari et al., ), and Lag-Llama (Rasul et al., 2023). We also consider 8 strong end-to-end supervised models, including multimodal ones like GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025a), and Time-VLM (Zhong et al., ), and unimodal ones like TSDiff (Kollovieh et al., 2023), CSDI (Tashiro et al., 2021), TimeGrad (Rasul et al., 2021a), and GRU MAF (Rasul et al., 2021b). The detailed information is provided in Appendix A.
@@ 4.3 Unimodal Forecasting
Models Aurora Sundial ROSE Timer TimesFM Chronos Time-MoE UniTS MOIRAI
(Ours) (2025) (2025) (2024) (2023) (2024) (2024) (2024) (2024)
ETT (Avg) 0.331 0.376 0.335 0.379 0.393 0.411 0.551 0.478 0.415 0.406 0.442 0.408 0.357 0.390 0.471 0.437 0.382 0.388
Weather 0.230 0.267 0.234 0.270 0.265 0.305 0.292 0.313 - - 0.288 0.309 0.256 0.289 0.275 0.298 0.260 0.275
Electricity 0.178 0.275 0.169 0.265 0.234 0.320 0.297 0.375 - - - - - - 0.198 0.291 0.188 0.273
Traffic 0.524 0.352 - - 0.588 0.412 0.613 0.407 - - 0.615 0.421 - - - - - -
Solar 0.203 0.289 0.221 0.252 0.505 0.549 0.771 0.604 0.500 0.397 0.393 0.319 0.411 0.428 0.845 0.669 0.714 0.704
PEMS08 0.563 0.552 - - 1.369 0.979 0.866 0.695 1.485 0.907 1.707 1.024 - - 1.253 0.879 - -
Wind 1.151 0.763 1.186 0.772 1.251 0.820 1.201 0.783 1.613 0.870 1.478 0.834 - - 1.425 0.848 1.299 0.795
1st Count 24 19 11 13 3 1 0 0 0 0 0 0 2 2 0 0 0 5
@@
1st Count 18 23 6 6 1 1 1 0 0 0 0 0 4 1 1 1 0 0
@@ 4.4 Model Analysis
Inference Scability. Adopting a generative probabilistic head, Aurora makes forecasts based on multiple sampling–see Algorithm 1. So that we study the scability of Prototype-Guided Flow Matching by exploring the correlations between the sampling number and forecasting performance in Figure 5. Specifically, experiments are conduct on ProbTS, where the average values of CRPS and NMAE across all datasets are reported. The results indicate that both CRPS and NMAE demonstrate a consistent improvement as the sampling number rises. They attain good performance when the sampling number reaches 100, showing obvious inference scability, and moderate efficiency–see Table 8.
@@ 5 Conclusion
In this work, we propose a highly capable multimodal time series foundation model, named Aurora. To sum up, Aurora adopts a merticulously-designed modality fusion process, which encodes, distills, then injects multimodal domain knowledge through a Modality-Guided Self-Attention into the modeling of temporal dynamics. To enhance the forecasting performance, Aurora utilizes multimodal representations to generate conditions and prototypes of future tokens, which contributes to a novel Prototype-Guided Flow Matching, thus supporting generative probabilistic forecasting. Comprehensive experiments on unimodal and multimodal forecasting tasks demonstrate that Aurora is a strong out-of-box tool for decision intelligence.

### v3

@@ Abstract
Cross-domain generalization is very important in Time Series Forecasting because similar historical information may lead to distinct future trends due to the domain-specific characteristics. Recent works focus on building unimodal time series foundation models and end-to-end multimodal supervised models. Since domain-specific knowledge is often contained in modalities like texts, the former lacks the explicit utilization of them, thus hindering the performance. The latter is tailored for end-to-end scenarios and does not support zero-shot inference for cross-domain scenarios. In this work, we introduce Aurora, a Multimodal Time Series Foundation Model, which supports multimodal inputs and zero-shot inference. Pretrained on Corss-domain Multimodal Time Series Corpus, Aurora can adaptively extract and focus on key domain knowledge contained in corrsponding text or image modalities, thus possessing strong Cross-domain generalization capability. Through tokenization, encoding, and distillation, Aurora can extract multimodal domain knowledge as guidance and then utilizes a Modality-Guided Multi-head Self-Attention to inject them into the modeling of temporal representations. In the decoding phase, the multimodal representations are used to generate the conditions and prototypes of future tokens, contributing to a novel Prototype-Guided Flow Matching for generative probabilistic forecasting. Comprehensive experiments on 5 well-recognized benchmarks, including TimeMMD, TSFM-Bench, ProbTS, TFB, and EPF, demonstrate the consistent state-of-the-art performance of Aurora on both unimodal and multimodal scenarios.
@@ 1 Introduction
Time series forecasting has gained sustained attention for decades of years due to its significant values in multiple domains, including economy, transportation, meteorology, and public health. In recent years, the key pivot comes with the surge of deep learning, which brings the boom of merticulously-designed deep forecasting models (Cirstea et al., 2022; Nie et al., 2023; Qiu et al., 2025d; Wu et al., 2025d). Through learning the inherent dynamics within the raw data, deep learning models can outperform classic statistical methods (Box & Pierce, 1970; Mei et al., 2014) and obey the scaling law (Shi et al., 2024a; Yao et al., ). Due to the success, it also brings the most commonly-used forecasting paradigm, which utilizes the past information to infer how the series goes in the coming horizon. Although this paradigm contributes to impressive performance under the domain-specific scenarios, its effectiveness is suspicious when facing cross-domain inference, where similar historical information may lead to different futures due to domain differences.
As shown in Figure 1, current research of time series forecasting explores the cross-domain adaption problem in two main perspectives: 1) pre-training on cross-domain time series corpus for unimodal time series foundation models, which partially possess cross-domain generalization capablities;
@@
2) utilizing cross-modality information in training end-to-end multimodal supervised models, which effectively integrates domain knowledge in forecasting. For time series foundation models, the cross-domain generalization capabilities mainly come from the sensitivity to subtle differences in historical information from different domains. Some of them (Shi et al., 2024b; Liu et al., 2025e) are pretrained on trillion-scale corpus with hevay backbones, thus possessing certain cross-domain adaption capabilities. Others (Wang et al., 2025c; b; Wu et al., 2025a) have specific structures, which excels at capturing cross-domain features. However, their capabilites come from single time modality and lack explict domain knowledge guidance, thus hindering the performance. For end-to-end multimodal supervised models (Jin et al., 2024; Liu et al., 2025a), though they consider the multimodal knowledge to enhance the domain-specific forecasting, they lack the ability to support zero-shot forecasting in cross-domain scenarios. In our view, the aurora of next-generation time series foundation model lies in pretraining a cross-modality model on cross-domain time series corpus, which can utilize the domain knowledge within modalities and serve as a versatile out-of-the-box forecaster in complex scenarios.
@@
  • Experimentally, Aurora achieves state-of-the-art performance on 5 well-recognized benchmarks, including datasets from TimeMMD (Liu et al., 2024b), TSFM-Bench (Li et al., 2025c), ProbTS (Zhang et al., 2024), TFB (Qiu et al., 2024), and EPF (Olivares et al., 2023), covering comprehensive scenarios, i.e., unimodal, multimodal, deterministic, and probabilistic, thus demonstrating a strong out-of-the-box tool of decision intelligence.
@@ 2.2 Time Series Foundation Models
To support cross-domain generalization, unimodal Time Series Foundation Models are widely studied. The majority of them adopt Tranformer-based architectures, which are pretrained on time series corpus of billion- or trillion- scale to obtain the strong generalization capabilities. Among them, Sundial (Liu et al., 2025e), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TTM (Ekambaram et al., 2024), Chronos (Ansari et al., ), UniTS (Gao et al., 2024), Timer (Liu et al., 2024e), and TimesFM (Das et al., 2024) demonstrate strong zero-shot forecasting performance on unimodal tasks, even outperforming those full-shot supervised models in many cases. Considering the forecasting paradigm, Sundial, MOIRAI, Chronos, and Lag-Llama (Rasul et al., 2023) also support probabilistic forecasting, which provides additional robustness and versatility for decision-making. Despite their endeavors to enhance cross-domain generalization capabilities, when historical series exhibit similarities, the forecasts they generate remain static. This lack of adaptability renders them unable to accommodate the diverse and changing real-world domains.
@@ 3.2.2 Prototype-Guided Flow Matching
Different from DDPM (Ho et al., 2020), which can be treated as an SDE solver to transform data from fixed Gaussian distributions to realistic target distributions, Flow Matching (Lipman et al., 2023) serves as a more intuitive and smooth ODE solver, which learns the Velocity Field between a random initial distribution and the target distribution. However, current methods (Liu et al., 2025e; Kollovieh et al., ) still set the initial distributions as Standard Gaussian, which neglects the capability of Flow Matching to work like a stochastic interpolant. Obviously, constructing a propriate prototype as the inital starting point can enhance the intuitiveness and stability of Flow Matching.
In the inference phase, the sampling process is a discretized integration process–see Algorithm 3.2.2. The gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,\mathbf{I})$ helps support probabilistic forecasting. Finally, we can obtain the forecasts $\hat{y}_{i}\in\mathbb{R}^{p^{time}}$ of the $i$-th future token. And the forecasts of the future horizon are $\hat{Y}=\text{Concat}\{\hat{y}_{i}\}\in\mathbb{R}^{F\times p^{time}}$.
@@ 4 Experiments
We make extensive experiments to evaluate the performance of Aurora. Specifically, we introduce the experimental settings in Section 4.1. In Section 4.2, we evaluate the zero-shot and few-shot performance of Aurora on multimodal forecasting scenarios. Considering the modal absence in the realistic world, we also evaluate the zero-shot performance of Aurora on unimodal forecasting scenarios–see Section 4.3, 4.4. To analyze the key components in Aurora, we also make detailed model analyses in Section 4.5. In summary–see Figure 4, our proposed Aurora achieves state-of-the-art performance in both unimodal and multimodal forecasting scenarios.
@@ 4.1 Experimental Settings
Cross-Domain Multimodal Time Series Corpus. We first collect a substantial number of open-source time series datasets across diverse domains, then generate the corresponding sample-wise textual descriptions using Large Language Model (Liu et al., 2024a), which simulates the downstream scenarios with domain-specific textual information.
Benchmarks. We evaluate both the multimodal forecasting and unimodal forecasting performance of Aurora on 5 benchmarks, including TimeMMD (Liu et al., 2024b), TSFM-Bench (Li et al., 2025c), ProbTS (Zhang et al., 2024), TFB (Qiu et al., 2024), and EPF (Wang et al., 2024b). Note that these benchmarking datasets are strictly excluded from the pretraining time series corpus.
Baselines. We compare Aurora with 11 well-known unimodal time series foundation models, including Sundial (Liu et al., 2025e), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TTM (Ekambaram et al., 2024), TimesFM (Das et al., 2024), Timer (Liu et al., 2024e), UniTS (Gao et al., 2024), Chronos (Ansari et al., ), and Lag-Llama (Rasul et al., 2023). We also consider multiple strong end-to-end supervised models, including multimodal ones like GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025a), and Time-VLM (Zhong et al., ), and unimodal ones like TimeXer (Wang et al., 2024b), PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024d) TSDiff (Kollovieh et al., 2023), CSDI (Tashiro et al., 2021), TimeGrad (Rasul et al., 2021a), and GRU MAF (Rasul et al., 2021b). The detailed information is provided in Appendix A.
@@ 4.3 Unimodal Forecasting
Models Aurora Sundial ROSE Timer TimesFM Chronos Time-MoE UniTS MOIRAI TTM
(Ours) (2025) (2025) (2024) (2023) (2024) (2024) (2024) (2024) (2024)
ETT (Avg) 0.331 0.376 0.335 0.379 0.393 0.411 0.551 0.478 0.415 0.406 0.442 0.408 0.357 0.390 0.471 0.437 0.382 0.388 0.441 0.430
Weather 0.230 0.267 0.234 0.270 0.265 0.305 0.292 0.313 - - 0.288 0.309 0.256 0.289 0.275 0.298 0.260 0.275 0.265 0.307
Electricity 0.178 0.275 0.169 0.265 0.234 0.320 0.297 0.375 - - - - - - 0.198 0.291 0.188 0.273 0.222 0.317
Traffic 0.524 0.352 - - 0.588 0.412 0.613 0.407 - - 0.615 0.421 - - - - - - 0.564 0.386
Solar 0.203 0.289 0.221 0.252 0.505 0.549 0.771 0.604 0.500 0.397 0.393 0.319 0.411 0.428 0.845 0.669 0.714 0.704 0.815 0.710
PEMS08 0.563 0.552 - - 1.369 0.979 0.866 0.695 1.485 0.907 1.707 1.024 - - 1.253 0.879 - - 1.730 1.066
Wind 1.151 0.763 1.186 0.772 1.251 0.820 1.201 0.783 1.613 0.870 1.478 0.834 - - 1.425 0.848 1.299 0.795 1.337 0.829
NYSE 0.528 0.526 0.880 0.642 - - 0.988 0.704 0.623 0.536 1.129 0.720 - - 1.220 0.820 - - - -
1st Count 27 21 11 13 3 1 0 0 1 2 0 0 2 2 0 0 0 5 0 0
@@
ILI 0.147 0.166 0.148 0.166 0.170 0.197 0.159 0.197 0.156 0.211 0.248 0.259 0.276 0.290 0.284 0.310 0.262 0.288
1st Count 19 24 8 8 1 1 2 1 0 0 0 0 4 1 1 1 0 0
4.4 Short-term Forecasting
Though some short-term forecasting settings are included in Section 4.2 and 4.3, we further evaluate Aurora on more unimodal short-term forecasting scenarios in this section, which are more in line with daily usage. Specfically, we conduct experiments on EPF (Wang et al., 2024b) and univariate datasets from TFB (Qiu et al., 2024). As shown in Table 4, Aurora outperforms most-advanced Foundation Models such as Sundial and VisionTS in most evaluations. Compared with full-shot supervised models like TimeXer, and iTransformer, Aurora also achieves competitive performance with them. Focusing on more scenarios, i.e., the 8,068 univariate datasets in TFB–see Figure 5, we report the mean MASE and msMAPE results, which indicate that Aurora also achieves state-of-the-art performance against zero-shot Foundation Models, and full-shot supervised models with versatile neural structures. All of the experiments demonstrate Aurora’s strong capability in short-term forecasting scenarios.
Table 4: Results of short-term zero-shot forecasting experiments on datasets from EPF. Lower MSE or MAE values indicate better predictions. (’-’) denotes datasets included in the model’s pretraining and therefore excluded from testing. Red: the best, Blue: the 2nd best.
Type [Uncaptioned image] Zero-shot Foundation Models [Uncaptioned image] Full-shot End-to-end Supversied Models
Models Aurora Sundial VisionTS ROSE MOIRAI TimeXer iTransformer PatchTST TimesNet
(Ours) (2025) (2025) (2025) (2024) (2024) (2024) (2023) (2023)
NP 0.288 0.312 0.256 0.277 0.510 0.461 0.666 0.536 0.660 0.538 0.238 0.268 0.265 0.300 0.267 0.284 0.250 0.289
PJM 0.084 0.183 0.088 0.189 0.251 0.366 0.311 0.402 0.330 0.423 0.088 0.188 0.097 0.197 0.106 0.209 0.097 0.195
BE 0.361 0.257 0.371 0.270 0.679 0.457 0.815 0.514 0.837 0.534 0.374 0.241 0.394 0.270 0.403 0.264 0.419 0.288
FR 0.387 0.206 0.392 0.207 0.625 0.393 0.746 0.447 0.751 0.454 0.381 0.211 0.439 0.233 0.411 0.220 0.431 0.234
DE 0.539 0.475 0.541 0.484 0.961 0.687 1.276 0.778 1.251 0.779 0.440 0.418 0.479 0.433 0.461 0.432 0.502 0.446
(a) Mean MASE results.
(b) Mean msMAPE results.
Figure 5: Mean MASE and msMAPE results of 8,068 univariate datasets in TFB. The full results can be found in Table 10 and 11 of Appendix B. Red: the best.
@@ 4.5 Model Analysis
Inference Scability. Adopting a generative probabilistic head, Aurora makes forecasts based on multiple sampling–see Algorithm 3.2.2. So that we study the scability of Prototype-Guided Flow Matching by exploring the correlations between the sampling number and forecasting performance in Figure 6. Specifically, experiments are conduct on ProbTS, where the average values of CRPS and NMAE across all datasets are reported. The results indicate that both CRPS and NMAE demonstrate a consistent improvement as the sampling number rises. They attain good performance when the sampling number reaches 100, showing obvious inference scability, and moderate efficiency–see Section A.7.
@@ 5 Conclusion
In this work, we introduce Aurora, the first multimodal time series foundation model. To sum up, Aurora adopts a merticulously-designed Modality-Guided Self-Attention to capture temporal dynamics, and a novel Prototype-Guided Flow Matching to enhance forecasting performance and supporting generative probabilistic modeling. Comprehensive experiments on unimodal and multimodal forecasting tasks, including 5 well-recognized benchmarks, demonstrate that Aurora is a strong out-of-the-box tool for decision intelligence.
Ethics Statement
Our work exclusively uses publicly available benchmark datasets that contain no personally identifiable information. The Cross-Domain Multimodal Time Series Corpus used to pretrain Aurora is also collected from public datasets, and integrated with LLM-generated textual descriptions, also containing no personally identifiable information. No human subjects are involved in this research.
Reproducibility statement
The performance of Aurora and datasets used in our work are real, and all experimental results can be reproduced. We have released our model code and checkpoints on Github and Huggingface.